# SPDFusion: An Infrared and Visible Image Fusion Network Based on a Non-Euclidean Representation of Riemannian Manifolds

### v2

@@ 4.3.4 Setting the Size of $\epsilon$
TABLE IV: Quantitative evaluation of object detection using the MSRS dataset.
Methods Year Person Car All Recall mAP@0.5 mAP@0.5:0.95
DenseFuse 2019 1.000 0.874 0.937 0.319 0.518 0.287
FusionGAN 2019 0.889 0.615 0.752 0.481 0.578 0.362
GANMcC 2021 0.609 0.883 0.746 0.705 0.784 0.424
RFN-Nest 2021 0.760 0.633 0.696 0.569 0.592 0.307
DeFusion 2022 0.882 0.696 0.789 0.690 0.825 0.473
FLFuse 2023 0.894 0.893 0.894 0.671 0.777 0.462
SemLA 2023 0.857 0.771 0.814 0.653 0.732 0.430
LRRNet 2023 0.850 0.634 0.742 0.671 0.797 0.469
DDFM 2023 0.932 0.801 0.867 0.629 0.810 0.483
SPDFusion Ours 1.000 0.879 0.939 0.695 0.810 0.491
@@ 4.4.1 Quantitative Comparisons
We performed an evaluation of the previously mentioned fusion networks using YOLOv7 [44]. The results were measured by the object detection metrics, namely: accuracy, recall, mean precision at 0.5 (AP50), and mean precision at 0.5:0.95 (AP50:95). As shown in Table IV, our method achieved the highest performance on three metrics, and secured the second-best performance on two detection metrics.
@@ 4.4.2 Qualitative Comparisons
In terms of visual effects, our method also has significant advantages, as shown in Fig. 12. In the examples of tradition methods, FusionGAN and RFN-Nest failed to detect the car, while DenseFuse, LRRNet and DDFM exhibited cases of redundant detection without accurately distinguishing the targets. When the picture shows night scenes with dim lighting, both GANMcC and FLFuse exhibit lower accuracy rates. Although DeFusion and SemLA have done well in the above aspects, false positives in regions with unclear edges have led to less than ideal detection results. In contrast, our SPDFusion maintained high detection accuracy in these challenging scenarios, preserving the prominent features and texture details of the targets, which further verifies the superiority of our approach.
@@ 5 Conclusion
Our attention module is designed based on the principle of convolution, using the covariance matrix as a weight matrix, which is then multiplied by the row vector matrix of the image. During this process, by utilizing the learned covariance matrix to adjust the weights of each pixel or pixel region, we can modulate the degree of ”attention” allocation to the image. This allocation mechanism reflects the intrinsic Riemannian geometrical structure of the image.
The results obtained on public datasets indicate that our approach preserves more detail features of visible light images and enhances the salient information of the infrared component, compared to existing fusion networks. Our fusion method has been applied to other computer vision tasks, such as object detection, and exhibited superior performance compared to other popular approaches as well.
In this work, we have only discussed one effective design of the SPD manifold in fusion tasks. It cannot be overemphasised that the research space for cross-modal fusion tasks based on manifold learning remains vastly expansive. In the future, we will continue to explore new methods for solving classical computer vision tasks using other manifolds.

### v3

@@ 4.3.4 The Impact of Different Manifold Constraints
Table 7: Quantitative evaluation of object detection using the MSRS dataset. The best three metrics are highlighted in bolditalic, bold and italic fonts, respectively.
Methods Person Car All mAP@0.5 mAP@0.5:0.95
DenseFuse (Li and Wu, 2018) 0.730 0.642 0.686 0.792 0.423
RFN-Nest (Li et al., 2021a) 0.619 0.611 0.615 0.750 0.482
DeFusion (Liang et al., 2022) 0.937 0.798 0.868 0.805 0.453
SemLA (Xie et al., 2023) 0.775 0.646 0.711 0.767 0.519
DDFM (Zhao et al., 2023b) 0.932 0.801 0.867 0.810 0.483
LRRNet (Li et al., 2023) 0.577 0.842 0.709 0.721 0.404
IRFS (Wang et al., 2023b) 0.773 0.799 0.786 0.819 0.482
CrossFuse (Li and Wu, 2024) 0.926 0.767 0.846 0.803 0.473
VDMUFusion (Shi et al., 2024) 0.750 0.514 0.632 0.760 0.439
DCINN (Wang et al., 2024) 0.611 0.622 0.617 0.761 0.456
DDBFusion (Zhang et al., 2025) 0.712 0.710 0.711 0.771 0.453
FusionBooster (Cheng et al., 2025b) 0.619 0.704 0.661 0.774 0.422
GIFNet (Cheng et al., 2025a) 0.512 0.740 0.626 0.693 0.388
SMLNet 0.939 0.997 0.968 0.822 0.545
Table 8: Quantitative evaluation of segmentation using the MSRS dataset. The best three metrics are highlighted in bolditalic, bold and italic fonts, respectively.
Methods Unl Car Per Bik mIOU
VI 97.89 74.75 35.31 62.59 67.63
IR 96.93 61.13 15.14 45.77 54.74
DenseFuse (Li and Wu, 2018) 97.82 73.91 24.70 65.37 65.45
RFN-Nest (Li et al., 2021a) 97.91 75.35 31.90 61.81 66.74
DeFusion (Liang et al., 2022) 97.85 74.18 21.84 64.70 64.64
SemLA (Xie et al., 2023) 97.84 73.87 32.96 59.95 66.16
DDFM (Zhao et al., 2023b) 97.85 72.90 33.37 61.33 66.36
LRRNet (Li et al., 2023) 97.86 73.57 40.06 59.21 67.68
IRFS (Wang et al., 2023b) 97.87 73.41 42.12 58.13 67.88
CrossFuse (Li and Wu, 2024) 97.86 73.70 33.00 62.07 66.66
VDMUFusion (Shi et al., 2024) 97.80 73.10 32.92 59.07 65.73
DCINN (Wang et al., 2024) 97.78 72.42 35.98 57.60 65.94
DDBFusion (Zhang et al., 2025) 97.85 73.06 33.22 63.15 66.82
FusionBooster (Cheng et al., 2025b) 97.87 72.78 34.38 65.30 67.58
GIFNet (Cheng et al., 2025a) 97.85 73.19 35.98 61.30 67.08
SMLNet 98.19 78.87 37.19 66.13 70.09
4.3.5 Comparison of Distributions under Different Learning Paradigms
In image fusion, effective fusion requires preserving complementary cross-modal interactions while respecting the intrinsic Riemannian geometry of the source data. To validate the geometric advantages of the manifold learning in SMLNet, we visualize original and dimensionality-reduced covariance matrices from MSRS samples using t-SNE. As shown in Fig. 13, each point corresponds to a row vector of the original covariance matrix, with Q1-Q4 representing statistical correlation blocks from four quadrants. To explicitly decouple modality-specific statistical signatures, we partition the covariance components into two geometrically meaningful groups: Q1/Q2 capturing modality-1-dominated covariance structures (infrared self-characteristics with cross-modal correlations) and Q3/Q4 representing modality-2-dominated covariance geometry (visible self-characteristics with complementary interactions). The t-SNE visualization result reveals: Traditional PCA forces linear projections onto Euclidean axes, flattening the curved SPD manifold structure and catastrophically compressing both modality-specific features (Q1/Q4) and cross-modal interactions (Q2/Q3). This distortion induces regional artifacts and thermal-textural misalignment (Rows 1 & 3). Conversely, our SPD transformations (BiMap layers) preserve intrinsic manifold geometry. Critically, t-SNE visualization confirms that our method uniformly blends modality-1-dominated and modality-2-dominated feature clusters on the SPD manifold. By performing dynamic Riemannian weighting within this geometrically consistent space, we optimally fuse complementary information while resolving inter-modal conflicts. For instance, in scenes like streetlights (Row 2), foliage (Row 4), and vehicles (Row 5), infrared thermal variations and visible texture gradients are jointly encoded through SPD-manifold-based cross-modal interactions. The result is fused imagery preserving thermal contrast while enhancing structural edges, well-aligned with human vision.
To quantitatively assess the fusion quality, we employ the following metrics: Silhouette Score (SS), Inter-Modal Distance Ratio (IMDR) and Cross-Modal Nearest Neighbor Ratio (CM-NNR). The SS is defined as:
$\text{SS}=\frac{1}{N}\sum_{i=1}^{N}\frac{b(i)-a(i)}{\max\{a(i),b(i)\}},$ (20)
where $a(i)$ is the average distance between sample $i$ and all other samples within the same modality, measuring intra-modal cohesion. $b(i)$ denotes the smallest average distance from sample $i$ to samples in any other modality, reflecting inter-modal separation.
IMDR evaluates fusion uniformity by dividing the mean inter-modal distance by the mean intra-modal distance:
$\text{IMDR}=\frac{D_{\text{inter}}}{D_{\text{intra}}},$ (21)
where $D_{\text{inter}}$ is the mean pairwise distance between samples from different modalities:
$D_{\text{inter}}=\frac{1}{|S_{\text{inter}}|}\sum_{(i,j)\in S_{\text{inter}}}d(x_{i},x_{j}),$ (22)
where $D_{\text{intra}}$ is the mean pairwise distance between samples within the same modality:
$D_{\text{intra}}=\frac{1}{|S_{\text{intra}}|}\sum_{(i,j)\in S_{\text{intra}}}d(x_{i},x_{j}),$ (23)
where $S_{\text{inter}}$ denotes the set of all cross-modal sample pairs and $S_{\text{intra}}$ represents the set of all intra-modal sample pairs. $d(x_{i},x_{j})$ is the distance metric (e.g., Euclidean distance) between samples $x_{i}$ and $x_{j}$.
CM-NNR quantifies mixing by averaging the fraction of neighbors from another modality for each point. For each sample $i$, the CM-NNR is calculated as:
$\text{CM-NNR}(i)=\frac{|\{j\in\text{NN}_{k}(i)|\text{label}(j)\neq\text{label}(i)\}|}{k},$ (24)
where $\text{NN}_{k}(i)$ is the set of k-nearest neighbors of sample $i$, $\text{label}(j)$ denotes the modality label of sample $j$, $k$ represents the number of nearest neighbors considered.
The global CM-NNR is computed as:
$\text{CM-NNR}=\frac{1}{N}\sum_{i=1}^{N}\text{CM-NNR}(i).$ (25)
As shown in Table 5, our SPD-manifold-based data points achieved the highest scores across all three metrics. The superior SS scores confirm enhanced feature compactness within modalities while maintaining clear separation between them. Optimal IMDR values demonstrate our method’s unique ability to balance cross-modal distances, while the highest CM-NNR indicates thorough modality mixing. These quantitative advantages directly validate that our SPD-manifold framework effectively eliminates heterogeneity while achieving tighter, more balanced fusion.
4.3.6 Noise Robustness in Riemannian vs Euclidean Networks
The presence of noise in real scene data is known to potentially distort feature representations and their statistical relationships, the robustness of multi-modal fusion methods heavily depends on their ability to maintain stable feature representations under noisy conditions. To further investigate this phenomenon, we conducted additional experiments.
At the input stage, we introduced controlled noise injection to artificially perturb the multi-modal data. As shown in Fig. 14, we systematically compared the fusion performance differences under varying noise levels. While Euclidean-based networks (especially ViT) demonstrate strong performance in low-noise scenarios due to the long-range modeling capability of attention mechanisms, they exhibit heightened sensitivity to cross-modal disturbances (e.g., discrete impulses from salt-and-pepper noise and continuous contamination from Gaussian noise), resulting in residual noise artifacts and blurring effects. In contrast, manifold-based networks model local statistical properties via covariance matrices. Since covariance captures second-order statistical relationships rather than raw pixel values, outliers are naturally suppressed during computation, enabling better preservation of edge structures and cross-modal features during denoising.
In Table 6, our manifold network delivers exceptional results in Gaussian noise removal, achieving near-state-of-the-art performance on both VIF and $Q^{AB/F}$ metrics. A noteworthy observation is that in salt-and-pepper noise scenarios, the SF metric exhibits non-linear correlation with fusion quality, attributable to the distinctive interference patterns of synthetic noise. Remarkably, our method achieves the optimal trade-off between information preservation and noise suppression, demonstrating the inherent stability of manifold-based feature representation.
@@ 4.4.1 Quantitative Comparisons
We performed an evaluation of the previously mentioned fusion networks using YOLOv7  (Wang et al., 2023a). The results were measured by the object detection metrics, namely: accuracy, mean precision at 0.5 (AP50) and mean precision at 0.5:0.95 (AP50:95). As shown in Table 7, our method achieved the highest performance on all five metrics.
@@ 4.4.2 Qualitative Comparisons
In terms of visual effects, our method also has significant advantages, as shown in Fig. 15. In the examples of tradition methods, SemLA failed to detect the car, while IRFS exhibited cases of redundant detection without accurately distinguishing the targets. Although CrossFuse have done well in the above aspects, false positives in regions with unclear edges have led to less than ideal detection results. In addition, FusionBooster has relatively limited capability in pedestrian detection. In contrast, our SMLNet maintained high detection accuracy in these challenging scenarios, preserving the prominent features and texture details of the targets, which further verifies the superiority of our approach.
4.5 Experiments in Semantic Segmentation
4.5.1 Quantitative Comparisons
To further validate the performance of our method on downstream vision tasks, we have conducted experiments on the MSRS segmentation dataset. During training, we used DeeplabV3+ (Chen et al., 2018) for network training, with the SGD optimizer and cross-entropy loss function for parameter updates. As shown in Table 8, we comprehensively measured the performance using the average accuracy and mIOU (Mean Intersection over Union) for 4 pixel-level category labels (car, person, bike and background).
SMLNet achieved the best segmentation effect on the background, car and bike categories, and obtained the highest mIOU. At the same time, the Person category also reached the 3rd level. This indicates that the manifold attention guides the intra-modal strongly correlated features to perform weight redistribution, ensuring the semantic consistency within the original image.
4.5.2 Qualitative Comparisons
Fig. 16 depicts a visual comparison between SMLNet and other competing methods. As demonstrated in the first two example rows, our method effectively highlights salient information from the infrared modality in challenging scenarios, while other methods exhibit severe misclassification. It is also noteworthy that, in daytime scenarios, the shape and position information of objects that are closely matched with the environmental colors can be accurately estimated, as seen in the last two rows of examples with the protrusion at the rear of the bicycles and the roof of the cars. In summary, these two typical examples vividly demonstrate the advantages of our method over other competitors.
Table 9: Efficiency comparison between SMLNet and 13 SOTA methods. The best three metrics are highlighted in bolditalic, bold and italic fonts, respectively.
Methods Params(MB) FLOPs(G)
DenseFuse 0.074 159.170
RFN-Nest 2.733 89.797
DeFusion 7.874 15.265
SemLA 7.278 16.613
DDFM 527.240 852.710
LRRNet 0.049 14.167
IRFS 0.242 78.655
CrossFuse 19.212 41.516
VDMUFusion 31.751 321.972
DCINN 2.097 17.026
DDBFusion 3.674 184.931
FusionBooster 0.555 56.407
GIFNet 0.613 39.814
SMLNet 0.051 3.652
4.6 Efficiency Comparison
In this section, we present the model complexity and computational cost of comparative methods. As shown in the Table 9, methods such as SemLA, CrossFuse, IRFS, and GIFNet adopt the classic ViT architecture, resulting in high computational costs. DDBFusion, DeFusion, and DCINN employ relatively complex decomposition modules, while DDFM and VDMUFusion rely on generative models, introducing significant computational overhead. DenseFuse, FusionBooster, LRRNet, and RFN-Nest utilize simpler network structures, offering advantages in computational efficiency. In contrast, our SMLNet, due to its lightweight network architecture and manifold dimensionality reduction design, demonstrates significant advantages compared to existing methods.
Figure 17: Comparative results of our fusion method under normal and extreme low-light conditions. The first and third rows display source images and their corresponding fusion results, while the second and fourth rows provide comparisons with other methods. In extreme low-light scenarios, the fusion quality is affected by error propagation from ill-conditioned covariance, leading to reduced contrast and subtle artifacts in dark regions. Under normal illumination, SMLNet effectively preserves critical details while integrating complementary cross-modal information.
4.7 Limitations in Extreme Scenarios Fusion
Though SMLNet demonstrates robust performance across most multi-modal fusion tasks, its effectiveness may degrade under extreme low-light conditions, such as nighttime scenes in the RoadScene dataset, while alternative methods maintain basic functionality. As shown in Fig. 17, in these challenging scenarios, our method occasionally struggles to clearly distinguish between tree and sky regions, while also exhibiting block artifacts and amplified noise in dark areas.
In such regions, critically weak signals cause local patches in both infrared and visible modalities to exhibit near-uniform intensity values. This lack of variation drives the eigenvalues of the computed covariance matrix $\mathbf{C}_{\mathbf{M}_{\text{ir}}\mathbf{M}_{\text{vi}}}$ toward zero, rendering it numerically singular. Although we employ SVD decomposition with a small perturbation term (Equation 6) to ensure positive definiteness, the extremely small eigenvalues still introduce numerical instability, affecting the subsequent dimensionality reduction in the BiMap layer and the reliability of SPD attention weights.
This issue is not a fundamental flaw of the method itself but arises because covariance matrix estimation relies on signal energy distribution. Under extremely low signal-to-noise ratio (SNR) conditions, noise dominates inter-modal relationships, degrading the accuracy of covariance-based feature mapping. In this scenario, global modeling approaches using second-order statistics show limited effectiveness, while methods relying on other principles may demonstrate relatively better performance. Consequently, future improvements should focus on enhancing covariance estimation robustness, optimizing eigenvalue conditioning, or exploring manifold-based fusion mechanisms adaptive to extreme SNR conditions.
@@ 5 Conclusion
Our attention module reformulates convolutional operations by treating the SPD matrix as a learnable weight kernel in eigenspace. This enables dynamic, geometry-aware feature modulation across pixel neighborhoods, directly responding to the intrinsic relationships within multi-modal data. By explicitly preserving the Riemannian structure during feature processing, our approach maintains critical high-frequency details and structural consistency in fused outputs, demonstrating superiority over Euclidean-space approaches.
Benchmark results on public datasets demonstrate that our approach not only retains richer visible-light details and enhances salient infrared features compared to existing fusion networks, but also achieves this with higher computational efficiency. Our fusion method has been applied to other computer vision tasks, such as object detection, and exhibited superior performance compared to other popular approaches as well.
In this work, we have only discussed one effective design of the SPD manifold in fusion tasks. It cannot be overemphasized that the research space for multi-modal fusion tasks based on manifold learning remains vastly expansive. In the future, we will continue to explore new methods for solving classical computer vision tasks using other manifolds.
Availability of Data and Materials
Information on access to the datasets supporting the conclusions of this article is included therein.
Competing Interests
The authors declare that they have no conflict of interest.
Acknowledgements.
This work was supported by the National Natural Science Foundation of China (62020106012, U1836218, 62106089, 62202205), the 111 Project of Ministry of Education of China (B12018), the Engineering and Physical Sciences Research Council (EPSRC) ( EP/V002856/1).